# Reflection Theory of Nichols Algebras over Coquasi-Hopf Algebras with Bijective Antipode

### v1

@@ 5.2 The semi-Cartan graph of a Nichols algebra
In this subsection, we develop reflections of Nichols algebras over arbitrary coquasi-Hopf algebra with bijective antipode. The reflection allows us to systematically relate different Nichols algebra realizations through a well-defined transformation procedure.
@@
Two tuples $M,M^{\prime}\in\mathcal{F}_{\theta}$ are called isomorphic, denoted $M\cong M^{\prime}$, if $M_{j}\cong M_{j}^{\prime}$ in $^{H}_{H}\mathcal{YD}$} for all j. The isomorphism class of $M\in\mathcal{F}_{\theta}$ is denote by $[M]$.
@@
$V_{j}=\begin{cases}M_{i}^{*},&\text{if }j=i,\\ (\mathrm{ad}\,M_{i})^{m_{ij}^{M}}(M_{j}),&\text{if }j\neq i.\end{cases}$
@@
By definition, $R_{i}(M)$ is defined, thus ${R_{i}(M)}_{i}\cong M_{i}^{*}$ is irreducible since $M_{i}$ is. For $j\neq i$, we observe that $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})=\bigoplus_{n=0}^{m_{ij}}\operatorname{ad}(M_{i})^{n}(M_{j})$. Consider the embedding $H\rightarrow\mathcal{A}(M_{i})$ and the projection $\mathcal{A}(M_{i})\rightarrow H$. The object $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ belongs to $^{H}_{H}\mathcal{YD}$}. Now suppose $0\neq P\subset\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ is a Yetter-Drinfeld submodule over $^{H}_{H}\mathcal{YD}$}. Let $\langle P\rangle$ be the $\mathcal{A}(M_{i})$-subcomodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ generated by $P$, defined explicitly as
@@
To establish this, we prove $\langle P\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}. To see this, for all $a\in\mathcal{A}(M_{i})$ and $x\in P$, by Lemma 2.6,
@@
Hence $\operatorname{ad}(a)(\langle f,x_{-1}\rangle x_{0})\in\langle P\rangle.$ This ensures the first axiom of a Yetter-Drinfeld module, and the third axiom holds automatically. Since $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ is generated by $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ as a Yetter-Drinfeld module and is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$} by Lemma 5.4 (2). Then $\langle P\rangle=\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ since $P\neq 0$. This forces $P=\operatorname{ad}(M_{i})^{m_{ij}}(M_{j}),$ establishing the irreducibility.
@@
(2) Suppose $M\cong N$ in $\mathcal{F}_{\theta}$, if $M$ admits the $i$-th reflection for some $i\in\mathbb{I}$, so is $N$. Furthermore, $R_{i}(M)\cong R_{i}(N)$ and $a_{ij}^{M}=a_{ij}^{N}$ for each $j\in\mathbb{I}$.
Recall by Remark 2.12, let $A$ be a Hopf algebra in $^{H}_{H}\mathcal{YD}$} with bijective antipode, we have such a braided monoidal isomorphism
@@
From now on, we always fix a tuple $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$, with each component $M_{i}$ is irreducible for $i\in\mathbb{I}$. By Corollary 4.3, we have such a braided tensor equivalence for each $i\in\mathbb{I}$.
@@
With above assumptions on $M$, and suppose $M$ admits the $i$-th reflection with $R_{i}(M)=(V_{1},\ldots,M_{i}^{*},\ldots,V_{\theta})$, then the following equalities hold.
@@
The next theorem gives a natural explanation of reflections of tuples of Yetter-Drinfeld modules. Although the proof strategy parallels the group case [Li and Liu, 2025b, Theorem 5.7], we include the details here for the sake of completeness and the reader’s convenience.
@@
With above assumptions on $M$, and suppose $M$ admits the $i$-th reflection, then there is an isomorphism of Hopf algebras in $^{H}_{H}\mathcal{YD}$}:
@@
Let $M\in\mathcal{F}_{\theta}$, with each $M_{i}$ is irreducible for $i\in I$. For $l\in\mathbb{N}_{0}$ and $i_{1},i_{2},...,i_{l}\in\mathbb{I}$.
$(1)$ We say $M$ admits the reflection sequence $(i_{1},i_{2},...,i_{l})$ if $l=0$ or $M$ amdits the $i_{1}$-th reflection and $R_{i_{1}}(M)$ satisfies the reflection sequence $(i_{2},i_{3},...,i_{l})$.
$(2)$ We say $M$ admits all reflection sequences if $M$ admits all reflection sequence $(i_{1},i_{2},...,i_{l})$ for all $l\in\mathbb{N}_{0}$ and $i_{1},i_{2},...,i_{l}\in\mathbb{I}$.
@@
$(CG1)$ For all $i\in I$, the map $r_{i}$ satisfies $r_{i}^{2}=\mathrm{id}_{\mathcal{X}}$.
$(CG2)$ For all $i\in I$, $X\in\mathcal{X}$, $A^{X}$ and $A^{r_{i}(X)}$ have the same $i$-th row.
@@
(1) For each $i\in\mathbb{I}$, suppose $M$ admits $i$-th reflection, then $R_{i}(M)$ admits $i$-th reflection. Furthermore, we have:
@@
Furthermore, we assume each $\mathcal{B}(P_{i})$ is finite-dimensional for $P\in\mathcal{X}$ and $i\in\mathbb{I}$, then
@@
is finite-dimensional. We have $\mathcal{B}(M)\simeq\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\#\mathcal{B}(M_{i})$ by Theorem 5.12. Further, $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}$ is generated by $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)$ by Theorem 5.5. Then we have $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i}$ generates $\mathcal{B}(M)$ and $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i}^{*}$ generates $\Omega_{i}(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})})\#\mathcal{B}(M_{i}^{*})$, and
@@
6 An example of affine Nichols algebra
In this section, we consider a rank three Nichols algebra $\mathcal{B}(M)$ of non-diagonal type in [Li and Liu, 2025b], which plays a central role in classification of finite-dimensional Nichols algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite abelian group and $\Phi$ is a $3$-cocycle on $G$. It was proved to be infinite-dimensional via different ways in [Huang et al., 2026,Li and Liu, 2025a,Li and Liu, 2025b]. Moreover, it gives rise to a semi-Cartan graph $\mathcal{G}(M)$. In this section, we prove $\mathcal{G}(M)$ is actually a Cartan graph and find a root system over $\mathcal{G}(M)$. Moreover, we show that $\mathcal{B}(M)$ is affine in the sense of Cuntz et al. (2019).
@@ 6.1 The Cartan graph of $\mathcal{B}(M)$
By [Huang et al., 2026, Lemma 3.5], we may give a complete list of irreducible Yetter-Drinfeld modules over $^{G}_{G}\mathcal{YD}^{\Phi}$}. In particular, restricting to those irreducible Yetter–Drinfeld modules that generate finite-dimensional Nichols algebras, the isomorphism classes are given by the set
@@
We now choose explicit representatives of $S$. For $1\leq i\leq 6$, let $M_{i}\in S$ be pairwise non-isomorphic simple modules such that $\operatorname{dim}(M_{i})=2$, and $\operatorname{deg}(M_{1})=h_{1}$, $\operatorname{deg}(M_{2})=h_{2}$, $\operatorname{deg}(M_{3})=h_{3}$, $\operatorname{deg}(M_{4})=h_{1}h_{2}$, $\operatorname{deg}(M_{5})=h_{1}h_{3}$, $\operatorname{deg}(M_{6})=h_{2}h_{3}$. One may refer to Li and Liu (2025b) for explicit expression of Yetter-Drinfeld structure, we omit here for simplicity.
The Nichols algebra $\mathcal{B}(M)$ which we concern is generated by $M_{1}\oplus M_{2}\oplus M_{3}$. Furthermore, we proved that:
@@
Here $\Phi_{A_{2}}=\pm\{\alpha_{2},\alpha_{3},\alpha_{2}+\alpha_{3}\}$, which corresponds to the root of Lie algebra $A_{2}$. This agreement is expected because the semi-Cartan graph is standard, and thus $\Delta^{M,\operatorname{re}}$ coincides with the set of real roots for the affine Lie algebra $A_{2}^{(1)}$. Unfortunately, $\Delta^{M,\operatorname{re}}$ is a infinite set, by definition $\mathcal{G}(M)$ is not a finite semi-Cartan graph.
@@
Then we say that the pair $(\mathcal{G},(R^{X})_{X\in\mathcal{X}})$ is a root system over $\mathcal{G}$. A root system over $\mathcal{G}$ is said to be reduced if for all $X\in\mathcal{X}$ and $\alpha\in R^{X}$ the roots $\alpha$ and $\alpha$ are the only rational multiples of $\alpha$ in $R^{X}$. A root system is said to be reduced if for all $X\in\mathcal{X}$, $R^{X}$ is finite.
@@
In [Cuntz, 2018,Cuntz et al., 2019], the authors showed that most of Nichols algebras over Hopf algebras can define a Tits cone, via its root system. In particular, they call a Nichols algebra affine if the corresponding Tits cone is a half plane. The following is their main observation:
@@
A crystallographic Tits arrangements consists of a pair $(\mathcal{A},T)$, where $\mathcal{A}$ is a set of linear hyperplanes in $\mathbb{R}^{r}$, and $T$ is a Tits cone in $\mathbb{R}^{r}$, satisfying some additional conditions. For explicit definition of crystallographic Tits arrangements, one may refer to [Cuntz et al., 2019, Definition 3.1].
We give a brief recall of calculating Tits cone when given a connected simply connected Cartan graph $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},(r_{i})_{i\in I},(A^{X})_{X\in\mathcal{X}})$ with a root system $R=(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}}$. Fix an object $A\in\mathcal{X}$. Let $V=\mathbb{R}^{r}$ (where $r=|I|$) and fix a linear isomorphism $\psi:\mathbb{Z}^{I}\to V^{*}$ sending the standard basis to a basis $\{\phi_{1},\dots,\phi_{r}\}$ of $V^{*}$. Define the set of roots $\mathcal{R}=\psi(\Delta^{A,\operatorname{re}})\subset V^{*}$.
@@
Theorem 6.7 guarantees that $(\mathcal{A},T)$ is a crystallographic Tits arrangement. The next proposition seems to be new, although those who are familiarly with affine Nichols algebras may regard it as a fact.
@@
We denote $r=|\mathbb{I}|$. Since $A$ is affine, it is automatically symmetrizable, then there exists a positive vector $v=(v_{1},v_{2},..,v_{r})\in\mathbb{Z}_{\geq 0}$ such that $v^{T}A=0$, that is
@@
Consider the affine plane $H_{1}=\{x\in V\mid\delta(x)=1\}$. The intersection of each chamber $K^{X}$ with $H_{1}$ is non‑empty because for any $x\in K^{X}$ we have $\delta(x)>0$, so $x/\delta(x)\in K^{X}\cap H_{1}$. Moreover, each chamber $K^{X}$ is a r‑dimensional open simplicial cone. The intersection $K^{X}$ with $H_{1}$ is a non‑empty (r1)(r-1)-dimensional simplex. In the theory of affine reflection groups, $\Delta:=K^{X}\cap H_{1}$ is known as a fundamental alcove. A classical result [Humphreys, 1990, Chapter 4] states that the affine Weyl group $W=W(A)$ acts simply transitively on the set of alcoves and that the closures of the alcoves tile the space $H_{1}$, i.e.
@@
The point $x/\delta(x)$ must lie in the closure of some simplex. Consequently, we can find finitely many vertices $p_{1},\dots,p_{n}$ of these triangles, where each vertex belongs to some chamber $K^{X_{i}}\cap H_{1}$ and non‑negative coefficients $t_{1},\dots,t_{n}$ with $\sum_{i=1}^{n}t_{i}=1$ such that
@@
Each $\delta(x)p_{i}$ belongs to $K^{X_{i}}$ because $K^{X_{i}}$ is a cone. Hence $x$ is a convex combination of points from $\bigcup_{X}K^{X}$, so $x\in T$. This proves $\delta^{+}\subseteq T$. Now we obtain $T=\delta^{+}$. Thus by philosophy of Theorem 6.7, the Tits cone corresponding to $(\mathcal{G},(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}})$ is a half plane. ∎
@@
(1) This means we realize an affine Nichols algebra over a coquasi-Hopf algebra, which broadens the realizing approaches of affine Nichols algebras.
(2) We hope to prove that a semi-Cartan graph induced by Nichols algebra over $^{H}_{H}\mathcal{YD}$} is actually a Cartan graph, which will endow them more beautiful properties.

### v2

@@ 5.2 The semi-Cartan graph of a Nichols algebra
In this subsection, we develop reflections of Nichols algebras over arbitrary coquasi-Hopf algebras with bijective antipode. The reflection allows us to systematically relate different Nichols algebra realizations through a well-defined transformation procedure.
@@
Two tuples $M,M^{\prime}\in\mathcal{F}_{\theta}$ are called isomorphic, denoted $M\cong M^{\prime}$, if $M_{j}\cong M_{j}^{\prime}$ in $^{H}_{H}\mathcal{YD}$} for all j. The isomorphism class of $M\in\mathcal{F}_{\theta}$ is denoted by $[M]$.
@@
$V_{j}=\begin{cases}M_{i}^{*},&\text{if }j=i,\\ \mathrm{ad}(M_{i})^{m_{ij}^{M}}(M_{j}),&\text{if }j\neq i.\end{cases}$
@@
By definition, $R_{i}(M)$ is defined, thus ${R_{i}(M)}_{i}\cong M_{i}^{*}$ is irreducible since $M_{i}$ is. For $j\neq i$, we observe that $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})=\bigoplus_{n=0}^{m_{ij}}\operatorname{ad}(M_{i})^{n}(M_{j})$. Consider the embedding $H\rightarrow\mathcal{A}(M_{i})$ and the projection $\mathcal{A}(M_{i})\rightarrow H$. The object $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ belongs to $^{H}_{H}\mathcal{YD}$}. Moreover, it is obvious that $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ is generated by $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ as a $\mathcal{A}(M_{i})$-comodule. Now suppose $0\neq P\subset\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ is a Yetter-Drinfeld submodule in $^{H}_{H}\mathcal{YD}$}. Let $\langle P\rangle$ be the $\mathcal{A}(M_{i})$-subcomodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ generated by $P$, defined explicitly as
@@
We want to prove $\langle P\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}. To see this, for all $a\in\mathcal{A}(M_{i})$ and $x\in P$, by Lemma 2.6,
@@
Hence $\operatorname{ad}(a)(\langle f,x_{-1}\rangle x_{0})\in\langle P\rangle.$ This ensures the first axiom of a Yetter-Drinfeld module, and the third axiom holds automatically. Therefore $\langle P\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}.
Since $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$} by Lemma 5.4 (2), and $P\neq 0$, we have $\langle P\rangle=\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$. It is direct to see that $\langle P\rangle=\bigoplus_{i=0}^{m_{ij}}\langle P\rangle\cap\mathcal{B}^{i}(M_{i}\oplus M_{j})$. Then
$P=\langle P\rangle\cap\mathcal{B}^{m_{ij}}(M_{i}\oplus M_{j})=\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})\cap\mathcal{B}^{m_{ij}}(M_{i}\oplus M_{j})=\operatorname{ad}(M_{i})^{m_{ij}}(M_{j}).$
This establishes the irreducibility. ∎
@@
(2) Suppose $M\cong N$ in $\mathcal{F}_{\theta}$, if $M$ admits the $i$-th reflection for some $i\in\mathbb{I}$, so does $N$. Furthermore, $R_{i}(M)\cong R_{i}(N)$ and $a_{ij}^{M}=a_{ij}^{N}$ for each $j\in\mathbb{I}$.
Recall from Remark 2.12, let $A$ be a Hopf algebra in $^{H}_{H}\mathcal{YD}$} with bijective antipode, we have such a braided monoidal isomorphism
@@
From now on, we always fix a tuple $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$, with each component $M_{i}$ being irreducible for $i\in\mathbb{I}$. By Corollary 4.3, we have such a braided tensor equivalence for each $i\in\mathbb{I}$.
@@
Under the above assumptions on $M$, and suppose $M$ admits the $i$-th reflection with $R_{i}(M)=(V_{1},\ldots,M_{i}^{*},\ldots,V_{\theta})$, then the following equalities hold.
@@
The next theorem gives a natural explanation of reflections of tuples of Yetter-Drinfeld modules. Although the proof strategy parallels that of group case [Li and Liu, 2025b, Theorem 5.7], we include the details here for the sake of completeness and the reader’s convenience.
@@
Under the above assumptions on $M$, and suppose $M$ admits the $i$-th reflection, then there is an isomorphism of Hopf algebras in $^{H}_{H}\mathcal{YD}$}:
@@
Let $M\in\mathcal{F}_{\theta}$, with each $M_{i}$ is irreducible for $i\in\mathbb{I}$. For $l\in\mathbb{N}_{0}$ and $i_{1},i_{2},...,i_{l}\in\mathbb{I}$.
$(1)$ We say $M$ admits the reflection sequence $(i_{1},i_{2},...,i_{l})$ if $l=0$ or $M$ admits the $i_{1}$-th reflection and $R_{i_{1}}(M)$ satisfies the reflection sequence $(i_{2},i_{3},...,i_{l})$.
$(2)$ We say $M$ admits all reflection sequences if $M$ admits reflection sequence $(i_{1},i_{2},...,i_{l})$ for all $l\in\mathbb{N}_{0}$ and $i_{1},i_{2},...,i_{l}\in\mathbb{I}$.
@@
$(CG1)$ For all $i\in\mathbb{I}$, the map $r_{i}$ satisfies $r_{i}^{2}=\mathrm{id}_{\mathcal{X}}$.
$(CG2)$ For all $i\in\mathbb{I}$, $X\in\mathcal{X}$, $A^{X}$ and $A^{r_{i}(X)}$ have the same $i$-th row.
@@
(1) For each $i\in\mathbb{I}$, suppose $M$ admits $i$-th reflection, then $R_{i}(M)$ admits the $i$-th reflection. Furthermore, we have:
@@
Then
@@
is finite-dimensional. We have $\mathcal{B}(M)\simeq\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\#\mathcal{B}(M_{i})$ by Theorem 5.12. Further, $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}$ is generated by $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)$ by Theorem 5.5. Then $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i}$ generates $\mathcal{B}(M)$ and $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i}^{*}$ generates $\Omega_{i}(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})})\#\mathcal{B}(M_{i}^{*})$, and
@@
6 Example of an affine Nichols algebra
In this section, we consider a rank three Nichols algebra $\mathcal{B}(M)$ of non-diagonal type in [Li and Liu, 2025b], which plays a central role in the classification of finite-dimensional Nichols algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite abelian group and $\Phi$ is a $3$-cocycle on $G$. It was proved to be infinite-dimensional in different ways in [Huang et al., 2026,Li and Liu, 2025a,Li and Liu, 2025b]. Moreover, it gives rise to a semi-Cartan graph $\mathcal{G}(M)$. In this section, we prove $\mathcal{G}(M)$ is actually a Cartan graph and find a root system over $\mathcal{G}(M)$. Moreover, we show that $\mathcal{B}(M)$ is affine in the sense of Cuntz et al. (2019).
@@ 6.1 The Cartan graph of $\mathcal{B}(M)$
By [Huang et al., 2026, Lemma 3.5], we may give a complete list of irreducible Yetter-Drinfeld modules over $^{G}_{G}\mathcal{YD}^{\Phi}$}. In particular, when restricting to those irreducible Yetter-Drinfeld modules that generate finite-dimensional Nichols algebras, we see that the isomorphism classes are given by the set
@@
We now choose explicit representatives of $S$. For $1\leq i\leq 6$, let $M_{i}\in S$ be pairwise non-isomorphic simple modules such that $\operatorname{dim}(M_{i})=2$, and $\operatorname{deg}(M_{1})=h_{1}$, $\operatorname{deg}(M_{2})=h_{2}$, $\operatorname{deg}(M_{3})=h_{3}$, $\operatorname{deg}(M_{4})=h_{1}h_{2}$, $\operatorname{deg}(M_{5})=h_{1}h_{3}$, $\operatorname{deg}(M_{6})=h_{2}h_{3}$. One may refer to Li and Liu (2025b) for the explicit expression of the Yetter-Drinfeld structure, we omit here for simplicity.
The Nichols algebra $\mathcal{B}(M)$ which we are concerned with is generated by $M_{1}\oplus M_{2}\oplus M_{3}$. Furthermore, we proved that:
@@
Here $\Phi_{A_{2}}=\pm\{\alpha_{2},\alpha_{3},\alpha_{2}+\alpha_{3}\}$, which corresponds to the root of Lie algebra $A_{2}$. This agreement is expected because the semi-Cartan graph is standard, and thus $\Delta^{M,\operatorname{re}}$ coincides with the set of real roots for the affine Lie algebra $A_{2}^{(1)}$. Unfortunately, $\Delta^{M,\operatorname{re}}$ is an infinite set, by definition $\mathcal{G}(M)$ is not a finite semi-Cartan graph.
@@
Then we say that the pair $(\mathcal{G},(R^{X})_{X\in\mathcal{X}})$ is a root system over $\mathcal{G}$. A root system over $\mathcal{G}$ is said to be reduced if for all $X\in\mathcal{X}$ and $\alpha\in R^{X}$ the roots $\alpha$ and $\alpha$ are the only rational multiples of $\alpha$ in $R^{X}$. A root system is said to be finite if for all $X\in\mathcal{X}$, $R^{X}$ is finite.
@@
In [Cuntz, 2018,Cuntz et al., 2019], the authors showed that most Nichols algebras over Hopf algebras can define a Tits cone, via their root systems. In particular, they call a Nichols algebra affine if the corresponding Tits cone is a half plane. The following is their main observation:
@@
A crystallographic Tits arrangement consists of a pair $(\mathcal{A},T)$, where $\mathcal{A}$ is a set of linear hyperplanes in $\mathbb{R}^{r}$, and $T$ is a Tits cone in $\mathbb{R}^{r}$, satisfying some additional conditions. For the explicit definition of crystallographic Tits arrangements, one may refer to [Cuntz et al., 2019, Definition 3.1].
We briefly recall the calculation of the Tits cone when given a connected simply connected Cartan graph $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},(r_{i})_{i\in I},(A^{X})_{X\in\mathcal{X}})$ with a root system $R=(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}}$. Fix an object $A\in\mathcal{X}$. Let $V=\mathbb{R}^{r}$ (where $r=|I|$) and fix a linear isomorphism $\psi:\mathbb{Z}^{I}\to V^{*}$ sending the standard basis to a basis $\{\phi_{1},\dots,\phi_{r}\}$ of $V^{*}$. Define the set of roots $\mathcal{R}=\psi(\Delta^{A,\operatorname{re}})\subset V^{*}$.
@@
Theorem 6.7 guarantees that $(\mathcal{A},T)$ is a crystallographic Tits arrangement. The next proposition seems to be new, although those who are familiar with affine Nichols algebras may regard it as a known fact.
@@
We denote $r=|\mathbb{I}|$. Since $A$ is affine, it is automatically symmetrizable; thus there exists a positive vector $v=(v_{1},v_{2},..,v_{r})\in\mathbb{Z}^{r}_{\geq 0}$ such that $vw3e2qA=0$, that is
@@
Consider the affine plane $H_{1}=\{x\in V\mid\delta(x)=1\}$. The intersection of each chamber $K^{X}$ with $H_{1}$ is non‑empty because for any $x\in K^{X}$ we have $\delta(x)>0$, so $x/\delta(x)\in K^{X}\cap H_{1}$. Moreover, each chamber $K^{X}$ is a r‑dimensional open simplicial cone. The intersection of $K^{X}$ with $H_{1}$ is a non‑empty (r1)(r-1)-dimensional simplex. In the theory of affine reflection groups, $\Delta:=K^{X}\cap H_{1}$ is known as a fundamental alcove. A classical result in [Humphreys, 1990, Chapter 4] states that the affine Weyl group $W=W(A)$ acts simply transitively on the set of alcoves and that the closures of the alcoves tile the space $H_{1}$, i.e.
@@
The point $x/\delta(x)$ must lie in the closure of some simplex. Consequently, we can find finitely many vertices $p_{1},\dots,p_{n}$ of these simplices, where each vertex belongs to some chamber $K^{X_{i}}\cap H_{1}$ and non‑negative coefficients $t_{1},\dots,t_{n}$ with $\sum_{i=1}^{n}t_{i}=1$ such that
@@
Each $\delta(x)p_{i}$ belongs to $K^{X_{i}}$ because $K^{X_{i}}$ is a cone. Hence $x$ is a convex combination of points from $\bigcup_{X}K^{X}$, so $x\in T$. This proves $\delta^{+}\subseteq T$. Now we obtain $T=\delta^{+}$. Thus, by the philosophy of Theorem 6.7, the Tits cone corresponding to $(\mathcal{G},(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}})$ is a half plane. ∎
@@
(1) This means we realize an affine Nichols algebra over a coquasi-Hopf algebra, which broadens the approaches to realizing affine Nichols algebras.
(2) We hope to prove that the semi-Cartan graph induced by a Nichols algebra over $^{H}_{H}\mathcal{YD}$} is actually a Cartan graph, which would endow it with more beautiful properties.